THE MILLIMETER/SUBMILLIMETER SPECTRUM OF METHYLPHOSPHINE, CH_{3}PH_{2} (X̃^{1}A)

\underline{D. T. HALFEN}, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ 85721; D. J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington, KY 40506; and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ 85721.

The millimeter/submillimeter spectrum of methylphosphine, CH_{3}PH_{2} (X̃^{1}A), has been measured using direct absorption techniques. Previously, only the microwave spectrum had been recorded. This molecule was created by the reaction of gas-phase phosphorus and methane or Si(CH_{3})_{4} in the presence of argon carrier gas and an AC glow discharge. Several transitions have been recorded in the range 280 - 422 GHz in both the v = 0 and v = 1 states each with multiple asymmetry components ranging from K_{a} = 0 to 16. Several of the K_{a} components in the v = 0 state show A/E splittings, while others appear collapsed. The v = 1 state has multiple K_{a} components with A/E splittings and is currently being analyzed. The data for the v = 0 state has been fit with an asymmetric top Hamiltonian, including internal rotation interactions, and the spectroscopic constants have been determined. Methylphosphine is the third row analog of methylamine, a known interstellar molecule, and could be a potential interstellar species.